A 0.10 M solution of formic acid (HCOOH) is prepared in various solvents: (a) pure water, (b) 0.10 M NaHCOO solution, (c) 0.10 M Ca(HCOO)2 solution, and (d) 0.10 M NaCl. Compute the pH of each solution. The Ka of formic acid is 1.78*10^-4.


1
Expert's answer
2021-03-12T06:29:26-0500

a)"pH=1\/2(pK_a-log(c_0))=2.38"

b)"pH=pK_a+log\left(\frac{c(NaHCOO)}{c(HCOOH)}\right)=pK_a=3.75"

c) "pH=pK_a+log\left(\frac{2c(Ca(HCOO)_2)}{c(HCOOH)}\right)=pK_a+log(2)=4.05"

d)"pH=2.38"